# No Detail Left Behind: Revisiting Self-Retrieval for Fine-Grained Image Captioning

### v1

@@ 4.8 Experiment 6: Combining Self-Retrieval and CIDEr Optimization
CIDEr optimization
(rows 3, 8) improves faithfulness to GT captions compared to MLE pretraining (rows 1, 6): CIDEr improves +10.7% on COCO and +29.1% on HolisticCap. Importantly, CIDEr optimization on top of HolisticCap MLE pretraining provides +4.8% on TrueMatch #7 (rows 6, 8), while we see only a smaller +2.5% with COCO captions. Even without SR fine-tuning, the larger improvement with HolisticCap illustrates the effectiveness of VCB and importance of MLE initialization with captions containing fine-grained visual details.
Furthermore, compared to vanilla SR-L (rows 2, 7), CIDEr opt. (rows 3, 8) fairs worse on both RD100 and TrueMatch. This presents an opportunity to optimize both rewards simultaneously (Luo et al., 2018).
@@
Building upon the findings of previous sections, we push the boundaries of SR by explicitly adding CIDEr to our discriminative reward $R=$ SR $+\lambda\cdot$ CIDEr. We conduct an ablation over various $\lambda$ values (see Section C.4), and select $\lambda=0.5$ for our experiments.
@@
Recall, our best model of row 10 also achieves highest performance on TrueMatch as compared to several MLLMs that are orders of magnitude larger and trained on more data (see Table 1).
@@ 5 Generalization Experiments
In this section, we evaluate the generalization of our approach on ImageCoDe (Krojer et al., 2022), a benchmark for text-based image retrieval where image sets are created with 10 sequential video frames. Different from TrueMatch, ImageCoDe focuses on fine-grainedness found across video frames such as negation, occlusion, etc. We follow the same experimental setup of Dessì et al. (2023). Table 8 shows that SR fine-tuning of both the vision and text modules with our bag curriculum (SR-LV BagCurri) achieves state-of-the-art results. In fact, it outperforms Dessì et al. (2023)’s work on using SR-L after MLE pretraining on COCO or Conceptual Captions (a 30×30$\times$30 × larger dataset).
@@ 6 Conclusion
We investigated self-retrieval (SR) as a way to comprehensively improve current captioning systems on all three fronts: (1) data, (2) evaluation, and (3) training. (1) We proposed Visual Caption Boosting (VCB), a model agnostic framework to instill fine-grained visual information into generic image captioning datasets while remaining anchored in human annotations. (2) Furthermore, we also created TrueMatch, a benchmark comprised of highly similar image bags. SR evaluation with TrueMatch enabled fine-grained evaluation of captioning systems and unlike traditional captioning metrics encouraged diversity in generated captions. (3) We also adopted SR fine-tuning and showed that the rich MLE initialization provided by VCB is important. We uncovered a trade-off between caption faithfulness and retrieval performance plaguing current SR approaches and designed a training recipe to address it. Specifically, we employed a curriculum over the number of hard negatives to more optimally leverage the SR reward, and fine-tune both the visual and language components of the captioning system. We demonstrated that this training recipe and VCB circumvents the aforementioned trade-off. Our approach achieved significant performance improvements over vanilla SR (Dessì et al., 2023) across a dataset with 99 random distractors, and set the state-of-the-art on ImageCoDe (Krojer et al., 2022). On TrueMatch, our approach outperforms powerful MLLMs that are 1-2 orders of magnitude larger and trained on large datasets, while improving over vanilla SR by up to +15%.
@@ 7 Limitations and Future Scope
Self-retrieval (SR) is bottlenecked by its reliance on a scorer to identify and retrieve the image that best matches the generated text. We employ CLIP as a scorer, which has been shown to struggle with capturing extremely fine-grained and compositional visual details. However, the reliability of SR evaluation is set to improve steadily over time. Additionally, our bag creation process for TrueMatch requires manual oversight to ensure that the benchmark consists of high-quality bags.
As future work, exploring further applications of the SR fine-tuning presents a promising research avenue. As we show, MLE initialization plays an important role and this too deserves more investigation.
Acknowledgements.
This project was supported in part by funding from SERB SRG/2023/002544 and an Adobe Research gift. We thank Shivanshu Sharma for further assistance with compute. We thank Lakshmipathi Balaji, Haran SK Raajesh, Naren Akash RJ, and Vansh Agarwal for early discussions.
@@ A.1 Prompting Strategy
We concurrently use InstructBLIP (Dai et al., 2024) to produce dense visual descriptions for the given image. We find the following prompt performs the best.
@@
We then prompt Mistral-7B to expand the blended caption with the visual description generated by the MLLM. Notably, the LLM is prompted to preserve human annotated information when there is a conflict of information between BlendCap and Visual Caption. Only 3 few shot examples are shown for brevity.
@@
A.2 Token Statistics for Captions
We present the caption statistics for Visual Caption Boosting compared to original COCO captions in Table 9. While Visual Captions are longer than HolisticCap, they may contain erroneous details as shown in Figure 2. Both BlendCap and HolisticCap are generally more descriptive than COCO, containing more words. Figure 4 shows the histogram of the number of tokens, tokenized using the CLIP tokenizer, for COCO, BlendCap and HolisticCap.
@@ A.2 Token Statistics for Captions
Table 9: Statistics of different captions. The mean and standard deviation of the number of words and tokens (post CLIP tokenization) are reported. Note the 4×4$\times$4 × difference between original COCO and HolisticCap.
@@ B.1 Bag Creation and Curation
3:      Set of $N$ images each with $M$ captions: =$\mathcal{I}=\{(i_{1},c_{1}^{1},\ldots,c_{1}^{M}),\ldots,(i_{N},c_{N}^{1},\ldots,c_{N}^{M})\}$.
@@ B.3 BagCurri: Designing a Self-Retrieval Learning Curriculum with Bags
We design a curriculum that varies the bag sizes during training, gradually increasing the bag size with each epoch as shown in Figure 5. Since the reward is the cross entropy of matching the generated caption with the target image Dessì et al. (2023), mining multiple hard negatives yields a flatter softmax distribution leading to stronger learning signal. Hence, increasing the bag size introduces makes the task of SR more challenging. However, it is important to increase bag sizes gradually to ensure that the task is not too hard, and prevent a collapse of caption faithfulness (Section 4.6).
@@ C.1 Discussion on Optimization Strategies
Similar to (Pinto et al., 2023), we find both optimization strategies discussed in Section 4.1 have complementary strengths and weaknesses for the task of image captioning.
@@ C.3 Overview of Hyperparameters
We provide a brief overview of hyperparameters for different optimization strategies: MLE (Table 12) and Reinforce (Table 12). Notably, CIDEr optimization for HolisticCap is done with learning rate $10^{-7}$.
@@
C.4 Scaling CIDEr during SR optimization
The weighting of the rewards when jointly optimizing CIDEr and SR (Luo et al., 2018) determines the trade-off between faithfulness and discriminative ability of the captioning system. We MLE pretrain the captioner with HolisticCap and fine-tune both the LLM and CLIP components (SR-LV) with BagCurri. Table 13 presents an ablation over different values of $\lambda$ while scaling CIDEr in the joint reward $R=$ SR $+\lambda\cdot$ CIDEr. As expected, increasing $\lambda$ i.e. heavily weighing CIDEr leads to improved caption faithfulness. Interestingly, using a smaller $\lambda$ however doesn’t always improve TrueMatch scores, as $\lambda=0.7$ (row 5) outperforms $\lambda=0.1$ (row 2). We adopt $\lambda=0.5$ for our experiments (Section 4.8), as it results in the best performing model achieving an optimal balance between retrieval performance and caption faithfulness.
@@ C.4 Scaling CIDEr during SR optimization
C.5 Fine-tuning CLIP hurts Caption Faithfulness
Data MLE SR-L SR-V
COCO 107.8 104 97.2
BlendCap 80.2 77.3 72.4
HolisticCap 26.6 26.4 25.6
@@ C.6 CIDEr Struggles with Longer Descriptions
C.7 Investigating Caption Unfaithfulness in SR Fine-tuning

### v2

@@ 4.8 Experiment 6: Combining Self-Retrieval and CIDEr Optimization
CIDEr optimization
(rows 3, 8) improves faithfulness to GT captions compared to MLE pretraining (rows 1, 6): CIDEr improves +10.7% on COCO and +29.1% on HolisticCap. Importantly, CIDEr optimization on top of HolisticCap MLE pretraining provides +4.8% on TrueMatch #7 (rows 6, 8), while we see only a smaller +2.5% with COCO captions. Even without SR fine-tuning, the larger improvement with HolisticCap emphasizes the importance of MLE initialization with captions containing fine-grained visual details (VCB). Furthermore, compared to vanilla SR-L (rows 2, 7), CIDEr opt. (rows 3, 8) fairs worse on both RD100 and TrueMatch. This presents an opportunity to optimize both rewards simultaneously (Luo et al., 2018).
@@
Building upon the findings of previous sections, we push the boundaries of SR by explicitly adding CIDEr to our discriminative reward $R=\text{SR}+\lambda\cdot\text{CIDEr}$. We conduct an ablation over various $\lambda$ values (see Section C.5), and select $\lambda=0.5$ for our experiments.
@@
Table 7 row 10 is our best model that also achieves highest performance on TrueMatch as compared to several MLLMs that are orders of magnitude larger and trained on more data (see Table 1).
@@ 5 Generalization Experiments
In this section, we evaluate the generalization of our approach on ImageCoDe (Krojer et al., 2022), a benchmark for text-based image retrieval where image sets are created with 10 sequential video frames. Different from TrueMatch, ImageCoDe focuses on fine-grainedness found across video frames such as negation, occlusion, etc. We follow the same experimental setup adopted by Dessì et al. (2023). Table 8 shows that SR fine-tuning of both the vision and text modules with our bag curriculum (SR-LV BagCurri) achieves state-of-the-art results. In fact, it outperforms Dessì et al. (2023)’s work on using SR-L after MLE pretraining on COCO (row 2, +7.6%) or Conceptual Captions, a 30×30$\times$30 × larger dataset (row 3, +1.7%).
@@ 6 Conclusion
We investigated self-retrieval (SR) as a way to comprehensively improve current captioning systems on all three fronts: (1) data, (2) evaluation, and (3) training. (1) We proposed Visual Caption Boosting (VCB), a model agnostic framework to instill fine-grained visual information into generic image captioning datasets while remaining anchored in human annotations. (2) Furthermore, we created TrueMatch, a benchmark comprised of highly similar image bags. SR evaluation with TrueMatch enabled fine-grained evaluation of captioning systems and unlike traditional captioning metrics encouraged diversity in generated captions. (3) We adopted SR fine-tuning and showed that the rich MLE initialization provided by VCB is important. We uncovered a trade-off between caption faithfulness and retrieval performance plaguing current SR approaches and designed a training recipe to address it. Specifically, we employed a curriculum over the number of hard negatives to more optimally leverage the SR reward, and fine-tune both the visual and language components of the captioning system. We demonstrated that this training recipe and VCB circumvents the aforementioned trade-off. With the same model architecture, our approach achieved significant performance improvements over vanilla SR (Dessì et al., 2023) across a dataset with 99 random distractors, and set the state-of-the-art on ImageCoDe (Krojer et al., 2022). On TrueMatch, our approach outperformed powerful MLLMs that are 1-2 orders of magnitude larger and trained on large datasets, while improving over vanilla SR by up to +15%.
@@ 7 Limitations and Future Scope
Self-retrieval (SR) is heavily reliant on a scorer to identify and retrieve the image that best matches the generated text. We employ CLIP as a scorer, which has been shown to struggle with capturing extremely fine-grained and compositional visual details. However, it is fair to expect that the reliability of SR evaluation should improve steadily over time with better vision-language models.
Additionally, as Visual Caption Boosting uses LLMs and MLLMs for generating fine-grained descriptions, it is prone to hallucinations. However, as shown, our anchoring strategy mitigates this to a large extent.
As future work, exploring further applications of the SR fine-tuning presents a promising research avenue. MLE initialization with improved captions is also a recurring theme and played an important role in our work – this too deserves further investigation while training MLLMs.
Acknowledgments
This project was supported in part by funding from SERB SRG/2023/002544 and an Adobe Research gift. We thank Shivanshu Sharma for partial assistance with compute. We thank Lakshmipathi Balaji, Haran SK Raajesh, Naren Akash RJ, and Vansh Agarwal for assistance with the human study.
@@ A.1 Prompting Strategy
We concurrently use InstructBLIP (Dai et al., 2024) to produce dense visual descriptions for the given image. We find that the following simple prompt performs well.
@@
We then prompt Mistral-7B to expand the blended caption with the visual description generated by the MLLM. Notably, the LLM is prompted to preserve human annotated information when there is a conflict of information between BlendCap and Visual Caption. Only 3 “few-shot” examples are shown here for brevity.
@@
A.2 Human Study: Quantifying Hallucinations in VCB
We conduct a human study to investigate the extent of hallucinations in BlendCap and HolisticCap. We randomly sample 60 images from MSCOCO and for each image-caption pair, a human is asked to count the number of hallucinations present in the caption. Three types of hallucinations are considered: (1) objects, (2) attributes, and (3) egregious statements, i.e., a glaringly incorrect statement that is unambiguously false.
BlendCap
contains only 4 object and 3 attribute hallucinations. Out of the 7 hallucinations, 5 are due to incorrect COCO human annotations and only 2 (3.3%) are induced by the LLM. This demonstrates the reliability of BlendCap in combining human annotations into a denser caption.
HolisticCap
contains only 9 instances of hallucinations (4 object and 5 attribute) across 7 captions, compared to VisualCap’s 16 instances (across 12 captions). Additionally, VisualCap contained 3 egregious statements that were removed in HolisticCap. Furthermore, 7/60 captions in HolisticCap contain hallucinations, reflecting a 41.6% reduction compared to VisualCap’s 12/60 captions. This demonstrates VCB’s effectiveness in anchoring the MLLM descriptions (Visual Captions) into human annotations (BlendCap) to create HolisticCap.
A.3 Token Statistics for Captions
We present some caption statistics for Visual Caption Boosting compared to original COCO captions in Table 9. While Visual Captions are longer than HolisticCap, they may contain erroneous details and verbose language modelling priors as shown in Figure 2. Both BlendCap and HolisticCap are generally more descriptive than COCO, containing more words. Figure 4 shows the histogram of the number of tokens, tokenized using the CLIP tokenizer, for COCO, BlendCap and HolisticCap.
@@ A.3 Token Statistics for Captions
Table 9: Statistics of different captions. The mean and standard deviation of the number of words and tokens (post CLIP tokenization) are reported. Note the 4×4$\times$4 × difference between the original COCO and HolisticCap. Also note that the verbosity of Visual Captions is reduced (by about 10) after combining with BlendCap to create HolisticCap.
@@ B.1 Bag Creation and Curation
3:      Set of $N$ images, each with $M$ captions: =$\mathcal{I}=\{(i_{1},c_{1}^{1},\ldots,c_{1}^{M}),\ldots,(i_{N},c_{N}^{1},\ldots,c_{N}^{M})\}$.
@@ B.3 BagCurri: Designing a Self-Retrieval Learning Curriculum with Bags
We design a curriculum that varies the bag sizes during training, gradually increasing the bag size with each epoch as shown in Figure 5. Since the reward is the cross-entropy of matching the generated caption with the target image Dessì et al. (2023), mining multiple hard negatives yields a flatter softmax distribution leading to stronger learning signal. Hence, increasing the bag size introduces makes the task of SR more challenging. However, it is important to increase bag sizes gradually to ensure that the task is not too hard, and prevent a collapse of caption faithfulness (Section 4.6).
@@ C.1 Discussion on Optimization Strategies
Similar to (Pinto et al., 2023), we find that both optimization strategies, MLE and Reinforce, discussed in Section 4.1 have complementary strengths and weaknesses for the task of image captioning.
@@ C.3 Overview of Hyperparameters
We provide a brief overview of hyperparameters for different optimization stages: MLE (Table 12) and Reinforce (Table 12). Notably, CIDEr optimization for HolisticCap is done with learning rate $10^{-7}$.
@@
C.4 Fine-tuning CLIP hurts Caption Faithfulness
Data MLE SR-L SR-V
COCO 107.8 104 97.2
BlendCap 80.2 77.3 72.4
HolisticCap 26.6 26.4 25.6
C.5 Scaling CIDEr during SR optimization
@@
The weighting of the rewards when jointly optimizing CIDEr and SR (Luo et al., 2018) determines the trade-off between faithfulness and discriminative ability of the captioning system. We MLE pretrain the captioner with HolisticCap and fine-tune both the LLM and CLIP components (SR-LV) with BagCurri. Table 14 presents an ablation over different values of $\lambda$ while scaling CIDEr in the joint reward $R=\text{SR}+\lambda\cdot\text{CIDEr}$. As expected, increasing $\lambda$ i.e. heavily weighing CIDEr leads to improved caption faithfulness. Interestingly, using a smaller $\lambda$ however doesn’t always improve TrueMatch scores, as $\lambda=0.7$ (row 5) outperforms $\lambda=0.1$ (row 2). We adopt $\lambda=0.5$ for our experiments (Section 4.8), as it results in the best performing model achieving an optimal balance between retrieval performance and caption faithfulness.
@@ C.6 CIDEr Struggles with Longer Descriptions
C.7 CHAIR: Quantifying Object Hallucinations during Extended SR-L Fine-tuning
We generate captions for all the 10,000 images of MSCOCO test and validation set and compute object level hallucinations using CHAIR (Rohrbach et al., 2018). CHAIR compares model-generated captions with human annotations and provides 2 scores $\text{CHAIR}_{I}$ and $\text{CHAIR}_{S}$ to quantify object and sentence level hallucinations respectively:
$\text{CHAIR}_{S}=\frac{|\{\text{captions with hallucinated objects}\}|}{|\{\text{all captions}\}|},\quad\text{CHAIR}_{I}=\frac{|\{\text{hallucinated objects}\}|}{|\{\text{all objects mentioned}\}|}.$ (2)
Table 15: Evaluating object hallucinations. We use CHAIR to study the extent of object hallucinations induced by MLE training and SR-L fine-tuning.
Dataset Method CHAIRI CHAIRS Hallucinated Objects CIDEr
COCO MLE 11.66 8.38 1255 107.8
SR-L 9.79 7.03 1033 108.2
SR-L 100 epochs 9.63 7.33 1102 92.4
MLE training
encourages the captioner to generate generic phrases for different images that may not be aligned with the image content, resulting in increased object hallucinations.
SR-L fine-tuning
may struggle with attribute binding, but it generates captions that are more discriminant and specific to the image, resulting in a significant reduction in object hallucinations.
Extended fine-tuning with SR-L
for 100 epochs in addition to reducing caption faithfulness, also increases both the absolute count and percentage of hallucinated objects.
C.8 Investigating Caption Unfaithfulness in SR Fine-tuning
@@
Figure 10: Manual filtering of bags in TrueMatch. We manually remove bags that do not capture some aspect of fine-grained visual discrimination, even if they represent the same visual concept.